Jason had to submit five animation projects as per his syllabus in the final semesters. He scored 82, 85, 78, and 88 points out of 100 in four projects. What should be his score in the fifth project so that the average of all the projects is at least 85?
Found 2 solutions by jojo14344, smarty1994:
Answer by jojo14344(1512)

Okay, let "x" be the number or score that Jason needs for the fifth project so he can have an average of at leat 85. To compute,
82+85+78+88+x=333+x, right? These are all his scores for the 5 projects. In order to get the average, we divide this by 5 and should equal to 85 right amigo? To show,
(333+x)/5 = 85 -------------------- eqn 1
solving for "x" thereafter,
333+x=85*5
x=425-333
x= 92.This should be his final score on the last project, what do you think?
In doubt?
Go back eqn 1,
(333+92)/5 = 85, cool!
Thank you,
Jojo

Answer by smarty1994(10)

(82+85+78+88)=5t
Add everything in the parenthesis and divide each side by 5 to get t = 66.6

Let t = the fifth test grade.